A square garden has an area of 30.25 square feet. What is the perimeter?
Answer by stanbon(75887)

A square garden has an area of 30.25 square feet. What is the perimeter?
----------------
Area = side^2
30.25 = s^2
s = sqrt(30.25) ft
----
Perimeter = 4s
P = 4*sqrt(30.25) = 22 ft
----------------
Cheers,
Stan H.
==============